64. 4w-1/3w+6 – w-1/3 = w-1/w+2
w=2

Answer by Edwin McCravy(3626)

<4w-1/3w+6 – w-1/3 = w-1/w+2
w=2

You'll have to put parentheses around numerators and denominators
and repost this problem. Otherwise we can't tell where the numerators
and denominators begin and where they end.
Please write it something like this:
4w - (1)/(32+6) - w - 1/3 = w - 1/(w+2)
or
(4w - 1)/(3w) - (6 - w - 1)/3 = (w - 1)/w + 2
or
(4w - 1)/(2w - 6) - (w - 3)/3 = (w - 1)/(w + 2)
because we can't tell what you mean.
Sorry.